# multimotion estimation problem MOT multiobject tracking

### v3

@@ 7.2 Estimator Comparison
MVO is a direct extension of VO to the MEP and can therefore leverage significant developments in egomotion estimation. This includes different $SE\left(3\right)$ estimation techniques, such as the discrete, WNOA, and WNOJ motion models presented here. The associated estimators have advantages and disadvantages in MVO. The pose-only estimator is the most general of those discussed in this paper. Its lack of motion prior makes it simple, but it can estimate physically implausible motions (e.g., discrete jumps) and it is not as robust to occlusions. The pose-velocity and pose-velocity-acceleration estimators constrain the trajectory estimate with a WNOA and WNOJ prior, respectively. This makes estimation more complex but penalizes physically implausible motions and allows for more accurate extrapolation and interpolation through occlusions. The WNOA prior assumes a locally constant velocity that is appropriate for many real-world situations but does not accurately model motions with large changes in direction. The WNOJ prior assumes a locally constant acceleration that is applicable to more real-world situations, including objects with large changes in direction; however, it often does not sufficiently constrain its estimates to accurately extrapolate motions for long periods of time.
@@ 7.2.2 Occlusion Segment
The estimators perform differently in the presence of occlusions, as seen in the occlusion_2_unconstrained results (Section 6.4). MVO consistently estimates the motions of both the camera and the moving blocks through motion closure with all estimators; however, the newly calculated centroid of a motion after an occlusion does not always match that of the original motion. This discrepancy causes jumps in the trajectory, as the original centroid is projected forward to a different location on the object than previously estimated. This affects the calculation of the correction transform (Section 5.3.2) and is a major source of error in the estimates of the block tower, as it is often partially outside the view of the camera. The global and relative estimation errors for the different estimators are given in LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:relative:occlusion and show that the pose-velocity estimator was generally the most accurate. The pose-only estimator performed reasonably well for directly observed motions but did not perform well on occlusions (Figs. 11, LABEL:tab:evaluation:global:occlusion, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:relative:occlusion). This is because its linear extrapolation and interpolation are sensitive to estimation noise, which often increases as an object becomes occluded (Section 7.1). The egomotion estimate is also affected by the sliding block transitioning between being stationary and dynamic more than the other estimators because of the lack of motion prior. The pose-velocity estimator performed well for the entire segment and was the best at extrapolating occluded motions (Figs. 12, LABEL:tab:evaluation:global:occlusion, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:relative:occlusion). Its locally constant-velocity assumption reasonably models many motions and allows for gradual changes in direction when directly observed. When changes in motion are not observed (i.e., during occlusion), the extrapolation can be inaccurate and motion closure may fail. This is expected because a WNOA prior only allows for small changes in velocity between consecutive frames, so it distributes large changes in motion over longer periods of time. The pose-velocity-acceleration estimator was the best at interpolating occluded motions, but occasionally underconstrained its extrapolated estimates (Figs. 13, LABEL:tab:evaluation:global:occlusion, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:relative:occlusion). Its locally constant-acceleration assumption directly models changes in velocity, even during occlusion. This more expressive prior results in better interpolations after motion closure than the WNOA prior but is less robust when extrapolating motions. This is because even small nonzero accelerations are extrapolated to significant motions over a moderate-length occlusion, as evidenced by the high extrapolation error for the swinging block. All estimators had high rotational error after occlusion because motion closure does not correct the extrapolated rotation (Section 5.3). This results in similar maximum rotation errors when extrapolating and interpolating (LABEL:tab:evaluation:global:swinging, LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:global:kitti). Accurately recognizing the relative rotation of an object after an occlusion is difficult without modeling the appearance or structure of the object. This may be achieved by tracking the visual features of the object or by building a model of its structure.
@@ 8 Conclusion
This paper presents MVO as an extension of the traditional VO pipeline to address the MEP. It uses multilabeling techniques to simultaneously segment and estimate all rigid motions in a scene. A variety of motion-based trajectory states are explored, and their effect on the ability to estimate independent $SE\left(3\right)$ motions and track them through occlusions is discussed. A stereo implementation of MVO is evaluated both quantitatively and qualitatively using the OMD (Judd and Gammell, 2019a) and the KITTI Vision Benchmark Suite (Geiger et al., 2012). MVO achieves egomotion estimation accuracy comparable to similarly defined egomotion-only VO systems, and it outperforms an appearance-based multimotion system (Zhang et al., 2020b) in addressing the MEP. The pipeline operates directly on tracked 3D points and is agnostic to the type of sensor that generates them, so a variety of sensors and corresponding estimators can be employed in MVO (Judd, 2019). MVO also relies exclusively on motion-based estimation and tracking techniques. Future work will include extending MVO to other sensors, such as LiDAR and event cameras, parallelizing the batch estimation in order to run at real time, and introducing application-specific, appearance-based techniques to improve performance in specific environments.
@@ 9 Acknowledgments
We would like to thank Paul Amayo for providing his implementation of CORAL and the authors of VDO-SLAM for providing their results and discussing them. We would also like to thank the editorial board for considering this manuscript and our reviewers for their time and constructive criticism. This research was funded by UK Research and Innovation and EPSRC through Robotics and Artificial Intelligence for Nuclear (RAIN) [EP/R026084/1], ACE-OPS: From Autonomy to Cognitive assistance in Emergency OPerationS [EP/S030832/1], and the Autonomous Intelligent Machines and Systems (AIMS) Centre for Doctoral Training (CDT) [EP/S024050/1].
@@ Appendix A Index to Multimedia Extensions
1 Video
Results of the pose-only estimator on the swinging_4_unconstrained segment of the OMD
2 Video
Results of the pose-velocity estimator on the swinging_4_unconstrained segment of the OMD
3 Video
Results of the pose-velocity-acceleration estimator on the swinging_4_unconstrained segment of the OMD
4 Video
Results of the pose-only estimator on the occlusion_2_unconstrained segment of the OMD
5 Video
Results of the pose-velocity estimator on the occlusion_2_unconstrained segment of the OMD
6 Video
Results of the pose-velocity-acceleration estimator on the occlusion_2_unconstrained segment of the OMD
7 Video
Results of the pose-velocity estimator on the Drive 0005 segment of KITTI
@@ Appendix B Nonlinear Stereo Camera Model
where $u$ and $v$ are the horizontal and vertical left-image coordinates and $d$ is the horizontal disparity between the image coordinate in the left and right image. This projection function has the corresponding Jacobian,

### v4

@@ 7.2 Estimator Comparison
MVO is a direct extension of VO to the MEP and can therefore leverage significant developments in egomotion estimation. This includes different $SE\left(3\right)$ estimation techniques, such as the discrete, WNOA, and WNOJ motion models presented here. The associated estimators have advantages and disadvantages in MVO.
The pose-only estimator is the most general of those discussed in this paper. Its lack of motion prior makes it simple, but it can estimate physically implausible motions (e.g., discrete jumps) and it is not as robust to occlusions. The pose-velocity and pose-velocity-acceleration estimators constrain the trajectory estimate with a WNOA and WNOJ prior, respectively. This makes estimation more complex but penalizes physically implausible motions and allows for more accurate extrapolation and interpolation through occlusions. The WNOA prior assumes a locally constant velocity that is appropriate for many real-world situations but does not accurately model motions with large changes in direction. The WNOJ prior assumes a locally constant acceleration that is applicable to more real-world situations, including objects with large changes in direction; however, it often does not sufficiently constrain its estimates to accurately extrapolate motions for long periods of time.
@@ 7.2.2 Occlusion Segment
The estimators perform differently in the presence of occlusions, as seen in the occlusion_2_unconstrained results (Section 6.4). MVO consistently estimates the motions of both the camera and the moving blocks through motion closure with all estimators; however, the newly calculated centroid of a motion after an occlusion does not always match that of the original motion. This discrepancy causes jumps in the trajectory, as the original centroid is projected forward to a different location on the object than previously estimated. This affects the calculation of the correction transform (Section 5.3.2) and is a major source of error in the estimates of the block tower, as it is often partially outside the view of the camera. The global and relative estimation errors for the different estimators are given in LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:relative:occlusion and show that the pose-velocity estimator was generally the most accurate.
The pose-only estimator performed reasonably well for directly observed motions but did not perform well on occlusions (Figs. 11, LABEL:tab:evaluation:global:occlusion, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:relative:occlusion). This is because its linear extrapolation and interpolation are sensitive to estimation noise, which often increases as an object becomes occluded (Section 7.1). The egomotion estimate is also affected by the sliding block transitioning between being stationary and dynamic more than the other estimators because of the lack of motion prior.
The pose-velocity estimator performed well for the entire segment and was the best at extrapolating occluded motions (Figs. 12, LABEL:tab:evaluation:global:occlusion, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:relative:occlusion). Its locally constant-velocity assumption reasonably models many motions and allows for gradual changes in direction when directly observed. When changes in motion are not observed (i.e., during occlusion), the extrapolation can be inaccurate and motion closure may fail. This is expected because a WNOA prior only allows for small changes in velocity between consecutive frames, so it distributes large changes in motion over longer periods of time.
The pose-velocity-acceleration estimator was the best at interpolating occluded motions, but occasionally underconstrained its extrapolated estimates (Figs. 13, LABEL:tab:evaluation:global:occlusion, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:relative:occlusion). Its locally constant-acceleration assumption directly models changes in velocity, even during occlusion. This more expressive prior results in better interpolations after motion closure than the WNOA prior but is less robust when extrapolating motions. This is because even small nonzero accelerations are extrapolated to significant motions over a moderate-length occlusion, as evidenced by the high extrapolation error for the swinging block.
All estimators had high rotational error after occlusion because motion closure does not correct the extrapolated rotation (Section 5.3). This results in similar maximum rotation errors when extrapolating and interpolating (LABEL:tab:evaluation:global:swinging, LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:global:kitti). Accurately recognizing the relative rotation of an object after an occlusion is difficult without modeling the appearance or structure of the object. This may be achieved by tracking the visual features of the object or by building a model of its structure.
@@ 7.2.3 KITTI Segments
The egomotion estimation error for the different estimators and VDO-SLAM on the Drive 0005 and Odometry 03 segments is given in LABEL:tab:evaluation:global:kitti and LABEL:tab:evaluation:relative:kitti and is less than $5.6\%$ of path length for all MVO estimators. These errors represent a similar percentage of the ground-truth motion as in the toy problems (LABEL:tab:evaluation:global:kitti, LABEL:tab:evaluation:global:total_path_kitti, LABEL:tab:evaluation:relative:kitti, 14 and 15). The segment does not include ground-truth trajectory data for third-party motions, so it is not useful for comparing the estimators beyond egomotion accuracy.
The decision of which estimator to use in any given application can be dependent on the requirements and characteristics of the situation. Based on the performance demonstrated on these segments, the WNOA is likely the most reliable prior for the general MEP. The WNOJ prior proved most accurate for interpolation, and it could likely be combined with the WNOA estimator for improved performance in cases with significant occlusion of nonconstant motions.
@@ 7.3 MVO Comparison
MVO and VDO-SLAM demonstrate similar tracking performance, but MVO outperforms VDO-SLAM for metric estimation of both ego- and third-party motions for the swinging_4_unconstrained (Section 6.3) and Drive 0005 (Section 6.5). MVO achieves this without relying on learning-based preprocessing steps or other appearance-based information and can track motions through significant occlusions.
VDO-SLAM prioritizes estimating relative object motions in outdoor scenes with minimal rotation, such as autonomous driving scenarios, and is not designed for significant occlusion, so it is not evaluated on occlusion_2_unconstrained. It is most comparable to MVO in egomotion accuracy, especially for the Drive 0005 segment (LABEL:tab:evaluation:global:kitti and LABEL:tab:evaluation:relative:kitti), but this design focus limits its applicability to the general MEP.
Unlike VDO-SLAM, MVO is reliant on motion, so objects that temporarily move similarly may be given the same label, such as when the sliding block tower becomes stationary (Section 6.4). Extrapolating lost motions even when they are similar to existing motions allows temporarily similar motions to be resegmented once they diverge, providing a form of motion permanence. This allows the block tower to be successfully tracked through motion-based occlusions, but MVO still had higher egomotion errors in the occlusion_2_unconstrained segment than the swinging_4_unconstrained segment because the block tower’s transitions between dynamic and stationary motion created noise in the egomotion label. VDO-SLAM uses front-end instance segmentation to avoid the difficulty of distinguishing between similarly moving objects, instead identifying objects to track based on appearance. Incorporating similar appearance-based techniques into MVO could improve its performance when independent objects temporarily move similarly.
@@
MVO and VDO-SLAM qualitatively demonstrate a similar ability to track both the cyclist and van in the Drive 0005 segment. Both consistently segment and estimate the cyclist until failing to track the van at the same time in the segment, though metric comparisons of their third-party estimates are not possible for this dataset due to its lack of ground-truth data. VDO-SLAM achieves this by relying on separate preprocessing algorithms to perform instance segmentation and optical flow calculation. MVO has no such dependencies, segmenting the scene and estimating trajectories based on the bulk motion of tracked feature points. This means it is applicable to a variety of different environments without significant tuning or prior knowledge of object number, appearance, or motion.
@@ 8 Conclusion
This paper presents MVO as an extension of the traditional VO pipeline to address the MEP. It uses multilabeling techniques to simultaneously segment and estimate all rigid motions in a scene. A variety of motion-based trajectory states are explored, and their effect on the ability to estimate independent $SE\left(3\right)$ motions and track them through occlusions is discussed.
A stereo implementation of MVO is evaluated both quantitatively and qualitatively using the OMD (Judd and Gammell, 2019a) and the KITTI Vision Benchmark Suite (Geiger et al., 2012). MVO achieves egomotion estimation accuracy comparable to similarly defined egomotion-only VO systems, and it outperforms an appearance-based multimotion system (Zhang et al., 2020b) in addressing the MEP.
The pipeline operates directly on tracked 3D points and is agnostic to the type of sensor that generates them, so a variety of sensors and corresponding estimators can be employed in MVO (Judd, 2019). MVO also relies exclusively on motion-based estimation and tracking techniques. Future work will include extending MVO to other sensors, such as LiDAR and event cameras, parallelizing the batch estimation in order to run at real time, and introducing application-specific, appearance-based techniques to improve performance in specific environments.
@@ 9 Acknowledgments
We would like to thank Paul Amayo for providing his implementation of CORAL and the authors of VDO-SLAM for providing their results and discussing them. We would also like to thank the editorial board for considering this manuscript and our reviewers for their time and constructive criticism.
This research was funded by UK Research and Innovation and EPSRC through Robotics and Artificial Intelligence for Nuclear (RAIN) [EP/R026084/1], ACE-OPS: From Autonomy to Cognitive assistance in Emergency OPerationS [EP/S030832/1], and the Autonomous Intelligent Machines and Systems (AIMS) Centre for Doctoral Training (CDT) [EP/S024050/1].
@@ Appendix A Index to Multimedia Extensions
1 Video Results of the pose-only estimator on the swinging_4_unconstrained segment of the OMD
2 Video Results of the pose-velocity estimator on the swinging_4_unconstrained segment of the OMD
3 Video Results of the pose-velocity-acceleration estimator on the swinging_4_unconstrained segment of the OMD
4 Video Results of the pose-only estimator on the occlusion_2_unconstrained segment of the OMD
5 Video Results of the pose-velocity estimator on the occlusion_2_unconstrained segment of the OMD
6 Video Results of the pose-velocity-acceleration estimator on the occlusion_2_unconstrained segment of the OMD
7 Video Results of the pose-velocity estimator on the Drive 0005 segment of KITTI
@@ Appendix B Nonlinear Stereo Camera Model
where $u$ and $v$ are the horizontal and vertical left-image coordinates and $d$ is the horizontal disparity between the image coordinate in the left and right image.
This projection function has the corresponding Jacobian,